Carl is 3 years older than lyn In five years the sum of their ages is 67. How old are they now?
Answer by samdom4peace(3)

Let Carl's age = C
Let Lyn's age = L
Carl is 3 years older than Lyn
C = L + 3 ...eqn. (1)
In five years,
Carl's age = C + 5
Lyn's age = L + 5
The sum of their ages is 67
(C + 5) + (L + 5) = 67
C + 5 + L + 5 = 67
C + L + 10 = 67
C + L = 57 ...eqn. (2)
Subst. (L + 3) for C in eqn. (2)
L + 3 + L = 57
2L = 54
L = 27
Subst. 27 for L in eqn. (1)
C = 27 + 3
C = 30
Carl is 30 years old; Lyn is 27 years old